# Thermodynamic Properties and Magnetocaloric Effect in a Rotating 2DEG under the Sagnac Effect

### v1

@@ Abstract
This work investigates the impact of the Sagnac effect on the thermodynamic properties of a non-interacting two-dimensional electron gas (2DEG) in a rotating sample under the influence of a uniform magnetic field. We analyze how rotation and the distinction between effective mass ($m^{*}$) and gravitational mass ($m_{G}$) affect the energy levels and resulting thermodynamic properties. The results show that rotation modifies the energy levels, the application of a magnetic field leads to the formation of Landau levels further altered by rotation and gravitational mass, and thermodynamic quantities (internal energy, specific heat, free energy, entropy, magnetization, and magnetocaloric effect) exhibit a strong dependence on these parameters. In particular, the difference between $m^{*}$ and $m_{G}$ influences magnetization and the magnetocaloric effect, with negative rotations potentially inducing a cooling effect when these masses are distinct. We conclude that rotational effects and effective mass properties are crucial for understanding the thermodynamics of electronic systems under magnetic fields, with implications for thermal modulation in semiconductor materials.
@@ I Introduction
where $\omega_{c}\equiv qB/m^{*}$ is the cyclotron frequency. The energy spectrum above is impacted by the Sagnac effect represented by the mass term $m_{G}$. We remark that the fields $B$ and $\Omega$ are externally tunable parameters. The presence of rotation by itself breaks the degeneracy between states with opposite angular momentum. It introduces energy shifts depending on the rotation ${\Omega}$ [24], which were analyzed in Ref. [24, 25], with $m_{G}\equiv m^{*}$. We will consider a more appropriate form for the energy levels. We can rewrite the energy levels in Eq. (11) as
@@ III.2.5 Magnetization
$M(T,B)=T\,\left(\frac{\partial\ln Z}{\partial B}\right)_{B,\Omega},$ (29)
@@ III.2.6 Magnetocaloric Effect
which describes the temperature variation with the magnetic field under isoentropic conditions. From it, we have

### v2

# Impact of the Sagnac Effect on Thermodynamic and Magnetocaloric Properties of a Rotating Two-Dimensional Electron Gas

@@ Abstract
This work investigates the impact of the Sagnac effect on the thermodynamic properties of a non-interacting two-dimensional electron gas (2DEG) in a rotating sample under the influence of a uniform magnetic field. We derive an analytical expression for the energy spectrum using an effective Hamiltonian that incorporates inertial forces, and then apply canonical ensemble statistical mechanics to evaluate thermodynamic quantities. The results show that rotation modifies the energy levels, the application of a magnetic field leads to the formation of Landau levels further altered by rotation and gravitational mass, and thermodynamic quantities (internal energy, specific heat, free energy, entropy, magnetization, and magnetocaloric effect) exhibit a strong dependence on these parameters. In particular, the difference between effective mass $m^{*}$ and gravitational mass $m_{G}$ influences magnetization and the magnetocaloric effect, with negative rotations potentially inducing a cooling effect when these masses are distinct. We conclude that rotational effects and effective mass properties are crucial for understanding the thermodynamics of electronic systems under magnetic fields, with implications for thermal modulation in semiconductor materials.
@@ I Introduction
Among the quantum systems where rotational and electromagnetic effects become particularly relevant, the two-dimensional electron gas (2DEG) stands out as a model system with rich physics and technological potential[21]. A 2DEG is a system in which electrons are confined to move in only two spatial dimensions, typically formed at the interface of semiconductor heterostructures such as GaAs/AlGaAs, or in devices like metal-oxide-semiconductor field-effect transistors (MOSFETs)[22]. The high mobility of these systems enables the observation of quantum phenomena such as the Quantum Hall Effect[23], making them essential for the study of low-dimensional systems and strongly correlated electrons[24]. Under strong magnetic fields, these systems exhibit remarkable quantum properties, including conductance quantization and the formation of Landau levels[25]. They have shown significance in optoelectronic applications as well[26]. The development of semiconductor nanomaterials and the exploration of quantum effects at the nanoscale are driving significant advancements in high-performance devices such as efficient solar cells, color micro-LEDs, advanced sensors, and innovative 3D displays[27].
@@
where $\omega_{c}\equiv qB/m^{*}$ is the cyclotron frequency. The energy spectrum above is impacted by the Sagnac effect represented by the mass term $m_{G}$. We remark that the fields $B$ and $\Omega$ are externally tunable parameters. The presence of rotation by itself breaks the degeneracy between states with opposite angular momentum. It introduces energy shifts depending on the rotation ${\Omega}$ [31], which were analyzed in Ref. [31, 32], with $m_{G}\equiv m^{*}$. Positive rotation ($\Omega>0$) refers to counterclockwise motion, while negative rotation ($\Omega<0$) refers to clockwise motion, following the right-hand rule convention with the angular velocity vector pointing along the $+z$ direction. We can rewrite the energy levels in Eq. (11) as
@@ III.2.5 Magnetization
$M(T,B)=T\,\left(\frac{\partial\ln Z}{\partial B}\right)_{T,\Omega},$ (29)
@@ III.2.6 Magnetocaloric Effect
which describes the temperature variation with the magnetic field under isoentropic conditions. $B_{i}$ and $B_{f}$ denote the initial and final values of the magnetic field, respectively, over which the adiabatic process and the MCE are evaluated. From it, we have